Hello ….. A ball falls freely from rest on a perfectly elastic surface from a height of 3 m . Draw the v- t graph for it upto 5 seconds.

Hello …..
      A ball falls freely from rest on a perfectly elastic surface from a height of 3 m . Draw the v- t graph for it upto 5 seconds.

Grade:9

1 Answers

Saurabh Koranglekar
3 years ago
The balls falls on a perfectly elastic platform, hence it will rebounce back without undergoing any loss of energy.

Hence, using second law of motion
s = ut + 1/2gt^2
u = 0, s = 3, g = 10
Hence, t = sqrt(60) = 2sqrt(15) sec

Also,the velocity with which it will hit the platform would be
v = u + at
v = 10*2sqrt(15)
v =20sqrt(15)

Hence, the ball will fall on the platform at t = 2sqrt(15) sec with a velocity of 20sqrt(15) m/s.
Since, the acceleration (retardation) is constant, hence velocity-time graph will initially be a straight line of positive slope from (0,0) to (2sqrt(15),20sqrt(15)) and then it will straight line with negative slope from(2sqrt(15),20sqrt(15)) to(4sqrt(15),0). This pattern of the graph will keep on repeating as there is no loss of energy.